OpenStudy (anonymous):

help pleasee The slope of a line is –4 and its y-intercept is (0, 4). What is the equation of the line that is perpendicular to the first line and passes through (–8, 2)?

OpenStudy (anonymous):

@Jamierox4ev3r

OpenStudy (sheraz12345):

Have u tried it urselves now?

OpenStudy (anonymous):

slope 1/4 right?

OpenStudy (anonymous):

y=1/4x+4 ?

OpenStudy (sheraz12345):

U told the slope right

OpenStudy (anonymous):

or is y= 1/4x-2

OpenStudy (sheraz12345):

but the equation is not right

OpenStudy (anonymous):

im confuse

OpenStudy (anonymous):

so is this one y= 1/4x-2

OpenStudy (sheraz12345):

oK let me show u thruogh diagram

OpenStudy (sheraz12345):

U were right sorry the answer is y=1/4x+4

OpenStudy (anonymous):

can you help in another one?

OpenStudy (sheraz12345):

Tell me

OpenStudy (anonymous):

Two perpendicular lines intersect at the origin. If the slope of the first line is -1/2 what is the equation of the second line?

OpenStudy (sheraz12345):

y=2x i think so

OpenStudy (anonymous):

A. y=-1/2x b. y = –2x c. y=2x

OpenStudy (sheraz12345):

c :D

OpenStudy (anonymous):

thanks

OpenStudy (sheraz12345):

Did u get it @RobertoTB ?